in 2018, Jeremy and Celeste, who file a joint return, paid the following amounts for their daughter, Alyssa, to attend the University of Colorado during academic year 2018-2019. Alyssa was in her first year of college and attended full-time.Problem 9-53 (LO 9-3) In 2018, Jeremy and Celeste, who file a joint return, paid the following amounts for their daughter, Al
Answer #1

Qualified expense are amounts in 2018 for tuition fees and enrolment and attendance costs.

Tuition and fees (for fall semester 2018)

1950

Tuition and fees (for spring semester 2019)

1000

Books

600

payment for her tuition

800

Qualifying expense for AOTC

$ 4,350

Modified AGI as given = $ 168,000.

The expense is more than 4000. The computation shall be done as the income falls in the phaseout range of 160,000 to 180,000.

The maximum deduction is 2500. The taxpayer is eligible for that as 100% of 2000 + 25% of next 2000 = $ 2500.

AOTC that can be claimed = 2500/20000   * (180,000-168000) = $ 1500

Note that for tax year 2018, the credit for tuition fees, supplies and equipment for the first years of college is limited to $ 2500. The limit phases out with income. If income goes above $ 160,000 for married and jointly filing taxpayers, the phaseout begins. Above 180,000, the phase ends.

In the given solution, the phaseout range is 180,000 – 160,000 = $ 20,000. This has been taken as denominator.

Further, the proportionate computation has been made

kindly upvote